A body is projected with velocity of 20 m/s at 50 degrees to the horizontal.find he maximum height,time of flight and range of the projectile
1
Expert's answer
2019-08-08T10:57:31-0400

The maximum height


"H=\frac{v_0^2 \sin^2\theta}{2g}=\frac{20^2 \sin^2 50^{\circ}}{2\times 9.8}=12\:\rm{m}"

The time of flight

"t=\frac{2v_0\sin\theta}{g}=\frac{2\times 20\times \sin50^{\circ}}{9.8}=3.13\:\rm{s}"

The range of the projectile


"R=\frac{v_0^2\sin 2\theta}{g}=\frac{20^2\times \sin 100^{\circ}}{9.8}=40.2\:\rm{m}"